Rx Chloral hydrate elixir    Sig: i dose of 250 mg, po hs How many teaspoonfuls...

Rx

Chloral hydrate elixir

   Sig: i dose of 250 mg, po hs

How many teaspoonfuls of chloral hydrate elixir with a strength of gr viiss/5 mL should be administered to the patient and at what time of day?
Answer #1

ANSWER.

* Calculation of teaspoon drug required.

         Formula

           Volume(ml) = Desired dose(D) / Dose in hand(H) * Quantity(Q).

Here

     Desired dose = 250 mg.

    Dose in hand(H) = 500 mg (gr viiss = 7.5 grains = 500 mg).

       Quantity(Q) = 5 ml.

As per above formula

     Volume(ml) of drug required = D/H * Q

                                                   = 250/500 * 5

                                                   = 0.5 * 5

                                                   = 2.5 ml.

              1 teaspoon = 5 ml.

   So

       Teaspoon of drug required = 2.5 / 5

                                                  = 0.5 tsp ( half of a teaspoon)

         * hs is a medical abbreviation for take at bedtime.

         * So this drug need to administer at bedtime.